OpenStudy (anonymous):

Factor out the GCF from the polynomial..... 14y^7+12y^2 . The correct answer to this is 2y^2(7y^5+6) . Can someone explain the process to do this, please? thanks...

OpenStudy (cwrw238):

first find the GCF of 14 and 12 this is the highest number which will divide exactly into 12 and 14 you can do this by factoring each number: (prime factors} 12 = 2*2*3 14 = 2*7 check for matching numbers:- there is one matching number 2 2 is the required GCF are you OK with this explanation?

OpenStudy (anonymous):

where do the other numbers come into play?

OpenStudy (cwrw238):

now deal with the 2 terms n 'y': y^7 = y*y*y*y*y*y*y y^2 = y*y y^2 is the GCF

OpenStudy (cwrw238):

you don't need to worry about the other numbers when finding the GCF - only the ones that match

OpenStudy (cwrw238):

to complete the factoring process you divide the GCF ( 2y^2) into the 2 terms of the original expression 14y^7 / 2y^2 = 7y^5 + 12y^2 / 2y^2 = + 6 to give 2y^2(7y^5 + 6)

OpenStudy (cwrw238):

ok? r u familiar with division of terms containg powers eg y^3 / y^2 ? you subtract the powers so y^3 / y^2 = y^(3 - 2) = y^1 = y

OpenStudy (anonymous):

stilll kinda confused..where does the 6 come from

OpenStudy (cwrw238):

12 / 2 = 6 y^2 / y^2 = 1

OpenStudy (cwrw238):

or you can treat it this way ( this is sometimes called 'cancelling'):